OpenStudy (anonymous):

An object is thrown upward following a parabolic path represented by y=-x^2+2x+11.At the same time a bird dives to catch the object and follows a path represented by y=7-x.

OpenStudy (anonymous):

A. When will the object and the bird be the same height from the ground?What does this mean in the context of the problem? B. When does the object reach its maximum height? C.When does the object hit the ground?

OpenStudy (amistre64):

when does y=y?

OpenStudy (welshfella):

they will be same distance from the ground at the instant the bird catches the object

OpenStudy (anonymous):

I dont know

OpenStudy (welshfella):

the graph will look something like this

OpenStudy (amistre64):

y=-x^2+2x+11 and y=7-x when does y=y? when -x^2+2x+11 = 7-x

OpenStudy (amistre64):

its assumed you know how to solve a quadratic equation ...

OpenStudy (welshfella):



OpenStudy (welshfella):

the point of intersection shows where the bird catches the object

OpenStudy (anonymous):

Is the point of intersection (4,-1)

OpenStudy (welshfella):

h is the height (y) and the time is x - in Amistre's equation